OpenStudy (anonymous):

Without using a calculator, find the values of the following expressions. (Give the answers in surd form if necessary.) (a) sin270°=? (b) tan360°=? (c) cos180°=? how to calculate?

hartnn (hartnn):

which one you have doubt with ?

OpenStudy (anonymous):

all. i can get the answer by using the calculator, but i don't know how to calculate them without it.

OpenStudy (anonymous):

Can I right 270 as 360 - 90 ?

OpenStudy (anonymous):

@kryton1212 , please reply.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

So I can also right : \(\sin 270 = \sin ( 360 - 90 ) \) Correct?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok! Take your time

OpenStudy (anonymous):

Good , now we have : the expression of the form \(\sin (360 - \theta)\) You know the formula for sin(360 - theta) ?

OpenStudy (anonymous):

-sinθ

OpenStudy (anonymous):

Great work. So \(\sin (360 - 90 ) = ? \)

OpenStudy (anonymous):

-sin90 ?

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

EXCELLENT!

OpenStudy (anonymous):

Ok! Do you know what is the value of \(\sin 360 \) and \(\cos 360\) ?

OpenStudy (anonymous):

it should be tan360

OpenStudy (anonymous):

0 and -1 respectively

hartnn (hartnn):

since, 360 =0, if you know tan 0, that will also do.

OpenStudy (anonymous):

Yes, alomost. \(\cfrac{\sin 360}{\cos 360} = \tan 360\) But here I am talking for the values of sin 360 and cos 360.

OpenStudy (anonymous):

Sin 360 = 0 is right but cos 360 = 1 not -1

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cos180=-1

OpenStudy (anonymous):